What are the measures of two complementary angles if the differences in the measures of two angles is 12

Answer by longjonsilver(2297)

complementary angles add up to 90 degrees.

Let smaller angle = x
so larger angle = x+12

x + x+12 = 90
2x+12 = 90
2x = 78
x = 78/2
x = 39

so angles are 39 and 51.

jon.